how many hectares is a piece of land which is 576 meter long and 450 meter wide?
with solution
tnx s answer
1
Expert's answer
2012-08-07T08:40:41-0400
The hectare is a metric unit of area defined as 10,000 square meters (100 m by 100 m).
Let's find area of our piece of lande:
S=576*450=259200 square meters
So we have 259200/10000=25,92 hectares